A unit vector is represented by a i^ +b j+c k^ . If a =0.6 and b=0.8 find c

Grade:11

5 Answers

sonika p
6 years ago
since, a i^ + bj^ + ck^ is a unit vector, 
root(a2 +b2 +c2) =1
a2+b2+c2=1
0.62  + 0.82  + c2 =1
0.36 + 0.64+ c2 =1
1 +c2 =1
c2 =0
thus, c =0
Gitanjali Rout
5 years ago
since, a i^ + bj^ + ck^ is a unit vector, 
root(a2 +b2 +c2) =1
a2+b2+c2=1
0.62  + 0.82  + c2 =1
0.36 + 0.64+ c2 =1
1 +c2 =1
c2 =0
thus, c =0jk
Gitanjali Rout
5 years ago
root(a2 +b2 +c2) =1
a2+b2+c2=1
0.62  + 0.82  + c2 =1
0.36 + 0.64+ c2 =1
1 +c2 =1
c2 =0
thus, c =0 hju ..............................................................................................
Gitanjali Rout
5 years ago
since, a i^ + bj^ + ck^ is a unit vector, 
root(a2 +b2 +c2) =1
a2+b2+c2=1
0.62  + 0.82  + c2 =1
0.36 + 0.64+ c2 =1
1 +c2 =1
c2 =0
Gitanjali Rout
5 years ago
since, a i^ + bj^ + ck^ is a unit vector, 
root(a2 +b2 +c2) =1
a2+b2+c2=1
0.62  + 0.82  + c2 =1
0.36 + 0.64+ c2 =1
1 +c2 =1
c2 =0 ….